OpenStudy (anonymous):

If F(x) = f(xf(xf(x))), where f(1) = 3, f(3) = 7, f '(1) = 3, f '(3) = 4, and f '(7) = 5, find F '(1).

zepdrix (zepdrix):

I don't quite get how you did that @Loser66 D: We don't have f(f(f(x))). We have f(x*f(x*f(x))). Or you were letting g and h equal x f(...) or something?

zepdrix (zepdrix):

I wasn't implying an f^3. I was reading it like this: \[\Large\rm F(x)=f\left[x\cdot f(x\cdot f(x))\right]\]Unless it was a typo I guess :o Which would differentiate something like this.\[\Large\rm F(x)=f'\left[x\cdot f(x\cdot f(x))\right]\cdot \color{royalblue}{(x\cdot f(x\cdot f(x)))'}\]And further uhhhh.....

zepdrix (zepdrix):

\[\large\rm F(x)=f'\left[x f(x f(x))\right]\cdot \color{royalblue}{\{f(x f(x))+x f'(x f(x)) f(x)+x f'(x))\}}\]Product rule and stuff.... D: Oh boy.. what a mess.. Sunyyyy, where you at?

OpenStudy (anonymous):

i think you are right

OpenStudy (anonymous):

but i am not sure how can i deal with the derivative of x(fx)

myininaya (myininaya):

you plug in at this point

OpenStudy (anonymous):

i guess the answer is 155 5*(4(3+3)+7)

zepdrix (zepdrix):

Hmm I got 125, maybe I made a boo boo somwhere. Where is that 4 coming from? Isn't that supposed to be a 3?

zepdrix (zepdrix):

I dunno, I can't make sense of this anymore XD and the website is tweaking out on me.. not letting me paste

myininaya (myininaya):

lol i got 59

myininaya (myininaya):

maybe i made a mistake

myininaya (myininaya):

well f'(3) is 4

myininaya (myininaya):

oops forgot some ( ) on mine found one of my mistakes

myininaya (myininaya):

ok 155 is what i got too zep

OpenStudy (anonymous):

thank you for your guys, the answer 155 is correct

myininaya (myininaya):

that was probably one of the ugliest chain rule product combinations i have ever seen requested

OpenStudy (anonymous):

yes, but all my homewok are all like this kind of chain rules combiations so sad